OpenStudy (anonymous):

Given the function f(x) = 2(x + 8), find x if f(x) = 12 2 40 −2 14

OpenStudy (anonymous):

f(x)= 12=2(x+8) 6= x+8 x=-2

OpenStudy (anonymous):

2(12+8) 2(20) 40

OpenStudy (anonymous):

:S

OpenStudy (saifoo.khan):

@Jithinkv 's right.

OpenStudy (saifoo.khan):

@Swagged_Solo : If the question was like: f(x) = 2(x + 8), find x if f(12) = 0 Then we would've used your method.

OpenStudy (anonymous):

@saifoo.khan oh iSee now !